# SASG-DA: Sparse-Aware Semantic-Guided Diffusion Augmentation For Myoelectric Gesture Recognition

### v2

@@ 4.2 Semantic Representation Guidance
Faithfulness is a key factor in data augmentation tasks. Specifically, it requires that synthetic samples preserve the distinctive features of their target classes, thereby ensuring that the downstream classifier learns from accurate and class-consistent information [wang2025inversion]. However, in conventional class-conditioned diffusion models, label conditions alone often provide only coarse-grained guidance, lacking sufficient semantic richness to generate highly faithful samples.
To overcome this limitation, we propose using fine-grained semantic representations as conditioning inputs. These continuous, class-specific features serve as informative guidance signals, allowing the model to generate more semantically aligned and realistic samples. Specifically, semantic representations $f\in\mathcal{R}^{D_{c}}$ are extracted from a pretrained task-aware classifier and integrated into the diffusion process via cross-attention mechanisms. We further apply dropout to the semantic features during training to improve robustness and mitigate overfitting to the conditioning input. In addition, label information is retained and incorporated through additive conditioning. The overall training objective combining (2) and (3) is modified as follows:
$\begin{split}\mathcal{L}_{\textit{SRG}}=&\mathbb{E}_{x_{0},x,t,y,f}\Big[\left\|x_{0}-\hat{x}_{0}(x_{t},t,y,f;\theta)\right\|^{2}\\ &+\left\|\mathcal{FFT}(x_{0})-\mathcal{FFT}(\hat{x}_{0}(x_{t},t,y,f;\theta))\right\|^{2}\Big]\end{split}$ (4)
@@ 4.3 Gaussian Modeling Semantic Sampling
In class-conditioned diffusion frameworks, the quality and diversity of the generated samples depend heavily on the design of the conditioning signal. To enable flexible and diverse generation, we propose a simple yet effective Gaussian Modeling Semantic Sampling (GMSS) mechanism. Specifically, we approximate the semantic representation distribution of each class as a multivariate Gaussian, a widely adopted approach in representation learning, which facilitates diverse sampling while maintaining semantic consistency.
@@
The sampled feature $\tilde{f}_{k}$ serves as a continuous condition to guide the reverse process of the diffusion model, enabling the generation of semantically aligned yet diverse gesture samples within class $k$.
Compared to label-based conditioning, which restricts generation to discrete and often coarse class labels, GMSS provides a continuous, distribution-based conditioning space that enables infinite and controllable sampling within each class. By sampling from different regions of the semantic representation distribution, GMSS allows the diffusion model to cover regions beyond the original training set while preserving semantic faithfulness. This not only increases intra-class diversity but also expands the effective data manifold, thereby improving the generalization performance of classifiers. In this way, diversity arises not only from the inherent stochasticity of the diffusion process but also from the stochastic variation introduced in the semantic condition space.
@@ 4.4 Sparse-Aware Semantic Sampling
While GMSS allows class-consistent sampling in a continuous semantic representation space, the diffusion model inherently tends to generate samples in dense regions of the feature distribution, thereby failing to explore underrepresented or sparsely populated areas [sehwag2022generating]. To address this limitation, we propose a Sparse-Aware Semantic Sampling (SASS) strategy that explicitly guides the generation process toward sparse regions in the data distribution. SASS employs a global–local sampling strategy: at the global level, we identify candidate features that are likely to lie in sparse regions of the distribution; at the local level, we optimize these candidates using a potential function that jointly encourages sparsity and diversity.
@@
To systematically identify potential sparse candidates within the oversampled pools, we adopt a rarity score metric proposed by [han2024rarity]. Specifically, we define the original training samples of each class as the reference set $R_{k}$ and establish $k$-NN spheres in the reference set, where each sphere is centered at a reference sample and its neighbors are also reference samples. For each candidate sample, the rarity score is computed as the radius of the smallest $k$-NN sphere that contains it. Intuitively, candidates located in denser regions will fall into smaller-radius spheres, while those closer to sparse regions will yield higher rarity scores. In addition, the rarity scores of outliers are set to zero. By ranking all candidates according to their rarity scores, we select high-rarity candidates as a preliminary pool of potential sparse conditions $\tilde{C}_{k}$.
Figure 3: Sparse-Aware Semantic Sampling. Candidate sparse features (blue) are optimized in the modeled semantic space using two potential functions. (a) The sparsity potential applies a repulsive force between candidate and reference features (orange), pushing candidates toward sparse regions. (b) The diversity potential encourages mutual repulsion among candidates to spread out within the semantic space.
@@
Formally, for each selected candidate feature $\tilde{f}$, we define its local neighborhood as the set of reference samples within an $\epsilon$-radius in the semantic space. In this way, each neighborhood is centered on the candidate feature, while its neighbors are reference samples from the same class. To make this neighbor count differentiable to gradient-based optimization, we replace the hard count with a smooth approximation using a sigmoid function $\sigma$:
@@
By minimizing this potential through gradient-based updates to the candidate feature $\tilde{f}$, the optimized features are adaptively repelled from high-density clusters and pushed into sparse regions of the class manifold.
Algorithm 1 Sparse-aware Semantic Sampling
1:Reference set $R_{k}$, Gaussian Modeling semantic representation distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$, neighborhood radius $\epsilon$, iterations iteriter, learning rate $\eta$.
2:Optimized sparse and diverse semantic representation set $S_{k}$.
3:Generate candidate set $C_{k}$ by sampling $10\times B$ features from $\mathcal{N}(\mu_{k},\Sigma_{k})$.
4:Compute rarity score for each $\tilde{f}_{i}$ in $C_{k}$.
5:Rank $\tilde{f}_{i}$ by rarity score descending, select top $B$ to form $\tilde{C}_{k}$.
6:for iteration $t=1$ to iteriter do
7:  Compute sparsity potential $\Phi_{\text{sparsity}}$ by (9).
8:  Compute diversity potential $\Phi_{\text{diversity}}$ by (10).
9:  Combine total objective: $\Phi=\Phi_{\text{sparsity}}+\Phi_{\text{diversity}}.$
10:  Compute gradient: $\nabla_{\tilde{C}_{k}}\Phi$.
11:  Update: $\tilde{C}_{k}\leftarrow\tilde{C}_{k}-\eta\cdot\nabla_{\tilde{C}_{k}}\Phi$.
12:end for
13:Evaluate classifier confidence for each $\tilde{f}_{i}$ in $\tilde{C}_{k}$; keep only those above threshold.
14:Supplement with high-fidelity $\tilde{f}_{i}\sim\mathcal{N}(\mu_{k},0.5\Sigma_{k})$.
15:return $S_{k}$.
@@
Minimizing $\Phi_{diversity}$ encourages candidate samples to spread apart within the feature space, discouraging trivial duplication and enforcing local sample-level diversity.
@@
Before feeding the optimized sparse features into the diffusion model as conditional guidance, we further ensure the quality of these conditions through a confidence-based filtering step. Specifically, we evaluate each optimized feature’s confidence score using the task-aware classifier and retain only those exceeding a predefined lower threshold. To compensate for discarded low-confidence samples and maintain balanced class coverage, the semantic representation conditions are further supplemented with high-fidelity samples drawn from GMSS, parameterized by the distribution $\mathcal{N}(\mu_{k},0.5\Sigma_{k})$. This ensures that the final condition set combines sparsity, diversity, and discriminative reliability for more effective diffusion-based augmentation. The overall algorithm is shown in Algorithm 1. To ensure efficiency and consistency, all condition features are synthesized offline, organized by class, before initiating the reverse diffusion process.
@@ 5.1 Datasets
To evaluate the effectiveness of the proposed approach, we conduct data augmentation experiments on three public sEMG benchmark datasets, namely Ninapro DB2 [atzori2014electromyography], DB4 [pizzolato2017comparison], and DB7 [krasoulis2017improved], as summarized in Table 1. These datasets encompass a diverse range of gesture classes and subjects, providing a comprehensive testbed for evaluating generalization. Each dataset contains multi-channel sEMG signals acquired using sparse-density acquisition systems with a sampling rate of 2 kHz. During preprocessing, the raw sEMG signals of each subject are segmented via a sliding window of length 200 ms with a step size of 40 ms, and then standardized channel-wise. To ensure consistency with previous studies [wang2024transformer, al2024tcnn] and focus on gesture-related categories, the rest class is excluded during preprocessing. Following established protocols [wang2024transformer, al2024tcnn] for gesture recognition based on sEMG, we adopt a cross-trial paradigm in which the training and testing sets are divided based on the trial information, as mentioned in Table 1.
Table 2: Performance comparison with SOTAs in terms of ACC (%), Pre(%), Rec(%) and F1(%) on dataset DB7 using three different backbones. Results are averaged among three randomized splits. Best performances are highlighted in bold.
@@
Baseline 77.97 78.86 78.37 78.61 73.54 74.99 73.96 74.46 73.61 75.10 74.17 74.63
Jittering & Scaling 78.05 79.12 78.44 78.78 74.10 75.44 74.53 74.92 72.28 73.98 72.88 73.43
Upsample 80.71 81.42 81.17 81.29 74.67 75.95 75.00 75.47 79.92 81.64 80.50 81.07
FreqMask 77.99 79.02 78.62 78.82 75.59 77.11 76.10 76.60 78.87 80.06 79.47 79.76
RIM 79.04 79.75 79.49 79.62 76.12 77.28 76.55 76.91 76.93 78.18 77.48 77.83
Dominant Shuffle 79.41 80.22 79.80 80.01 75.21 77.03 75.67 76.34 78.92 80.37 79.46 79.91
Mixup 76.99 77.79 77.58 77.68 75.17 76.03 75.68 75.85 73.90 75.18 74.55 74.86
STAug 77.43 78.30 77.95 78.12 75.32 76.29 75.72 76.00 73.53 74.77 74.22 74.49
FreqMix 79.79 80.96 80.22 80.59 76.24 77.68 76.56 77.12 78.35 79.94 78.86 79.40
SimPSI 78.23 78.94 78.60 78.77 74.12 75.43 74.52 74.97 74.51 75.92 75.16 75.54
DiffMix 78.14 78.90 78.52 78.71 74.44 75.40 74.86 75.13 74.26 75.51 74.89 75.20
Minority 79.86 80.30 80.37 80.33 75.14 75.86 75.67 75.76 79.84 80.53 80.37 80.45
CADS 80.41 80.83 80.86 80.84 76.36 77.16 76.75 76.95 80.55 81.28 81.07 81.17
SASG-DA (Ours) 81.31 81.89 81.75 81.82 78.77 79.73 79.23 79.48 82.15 83.07 82.68 82.87
Table 3: Performance comparison with SOTAs in terms of ACC (%), Pre(%), Rec(%) and F1(%) on dataset DB4 using three different backbones. Results are averaged among three randomized splits. Best performances are highlighted in bold.
@@
Baseline 65.78 66.97 67.49 67.23 62.62 64.18 64.34 64.26 61.25 63.43 63.05 63.24
Jittering & Scaling 66.01 67.35 67.80 67.57 63.05 64.60 64.75 64.67 59.97 62.23 61.74 61.98
Upsample 67.62 69.23 69.29 69.26 63.37 65.16 65.23 65.19 68.44 70.40 70.14 70.27
FreqMask 66.67 68.08 68.35 68.21 64.50 66.31 66.23 66.27 67.08 68.99 68.82 68.90
RIM 67.18 68.31 68.90 68.60 64.47 65.75 66.21 65.98 65.08 66.58 66.77 66.67
Dominant Shuffle 67.21 68.41 68.96 68.68 64.22 66.01 66.06 66.03 65.89 67.89 67.69 67.79
Mixup 66.21 67.48 67.86 67.67 64.50 65.96 66.20 66.08 62.13 63.95 63.82 63.88
STAug 66.41 67.77 68.09 67.93 63.96 65.31 65.67 65.49 61.46 63.36 63.10 63.23
FreqMix 68.51 70.03 70.25 70.14 65.45 67.08 67.15 67.11 66.69 68.73 68.38 68.55
SimPSI 66.72 67.93 68.40 68.16 64.06 65.54 65.78 65.66 62.44 64.48 64.06 64.27
DiffMix 66.08 67.21 67.81 67.51 62.14 63.35 63.83 63.59 63.22 64.89 64.99 64.94
Minority 68.39 69.49 70.03 69.76 64.47 65.86 66.13 65.99 66.74 68.38 68.42 66.40
CADS 68.91 69.99 70.57 70.28 65.58 67.03 67.33 67.18 67.45 68.84 69.13 68.98
SASG-DA (Ours) 69.73 70.73 71.38 71.05 67.74 69.02 69.42 69.22 70.05 71.37 71.75 71.56
Table 4: Performance comparison with SOTAs in terms of ACC (%), Pre(%), Rec(%) and F1(%) on dataset DB2 using three different backbones. Results are averaged among three randomized splits. Best performances are highlighted in bold.

### v3

@@ 4.2 Semantic Representation Guidance
Faithfulness is a key factor in data augmentation tasks. Specifically, it requires that synthetic samples preserve the distinctive features of their target classes, thereby ensuring that the downstream classifier learns from accurate and class-consistent information [8]. However, in conventional class-conditioned diffusion models, label conditions alone often provide only coarse-grained guidance, lacking sufficient semantic richness to generate highly faithful samples.
To overcome this limitation, we propose using fine-grained semantic representations as conditioning inputs. These continuous, class-specific features serve as informative guidance signals, allowing the model to generate more semantically aligned and faithful samples. Specifically, semantic representations $f\in\mathcal{R}^{D_{c}}$ are extracted from a pretrained task-aware classifier and integrated into the diffusion process via cross-attention mechanisms. The pretrained classifier is trained on the same gesture recognition task and dataset as the diffusion model, ensuring that its extracted features are discriminative and task-relevant, thus providing meaningful semantic guidance for generation. We further apply dropout to the semantic features during training to improve robustness and mitigate overfitting to the conditioning input. In addition, label information is retained and incorporated through additive conditioning. The overall training objective combining (2) and (3) is as follows:
$\begin{split}\mathcal{L}_{\textit{SRG}}=&\mathbb{E}_{x_{0},x,t,y,f}\Big[\left\|x_{0}-\hat{x}_{0}(x_{t},t,y,f;\theta)\right\|^{2}\\ &+\left\|\mathcal{FFT}(x_{0})-\mathcal{FFT}(\hat{x}_{0}(x_{t},t,y,f;\theta))\right\|^{2}\Big].\end{split}$ (4)
@@ 4.3 Gaussian Modeling Semantic Sampling
In class-conditioned diffusion frameworks, the faithfulness and diversity of the generated samples depend heavily on the design of the conditioning signal. To enable flexible and diverse generation, we propose a simple yet effective Gaussian Modeling Semantic Sampling (GMSS) mechanism. Specifically, we approximate the semantic representation distribution of each class as a multivariate Gaussian, a widely adopted approach in representation learning, which facilitates diverse sampling while maintaining semantic consistency.
@@
The sampled feature $\tilde{f}_{k}$ is used as a continuous condition to guide the diffusion reverse process, generating semantically aligned yet diverse samples for class $k$.
Compared to discrete and coarse-grained label-based conditioning, GMSS provides a continuous, distribution-based conditioning space that enables flexible sampling within each class. By sampling from different regions of the semantic space, GMSS expands coverage beyond the original training set while preserving semantic faithfulness. This not only increases intra-class diversity but also expands the effective data manifold, thereby improving the generalization performance of classifiers. Thus, diversity arises not only from inherent diffusion stochasticity but also from the stochastic variation introduced in the semantic condition space.
@@ 4.4 Sparse-Aware Semantic Sampling
While GMSS allows class-consistent sampling in a continuous semantic representation space, the diffusion model inherently tends to generate samples in dense regions of the feature distribution, thereby failing to explore underrepresented or sparsely populated areas [24]. These dense regions correspond to well-represented and frequently observed samples in the training data, whereas the sparse regions contain rare or insufficiently represented cases that the model has rarely seen but are crucial for improving downstream performance. To address this limitation, we propose a Sparse-Aware Semantic Sampling (SASS) strategy that explicitly guides the generation process toward sparse regions in the data distribution. SASS employs a global–local sampling strategy: at the global level, we identify candidate features that are likely to lie in sparse regions of the distribution; at the local level, we optimize these candidates using a potential function that jointly encourages sparsity and diversity.
@@
To identify potential sparse candidates from the oversampled pools, we adopt the rarity score proposed in [28]. For each class, the original training samples $R_{k}$ serve as the reference set to construct $k$-NN spheres, where both the centers and their neighbors are reference samples. The rarity score of a candidate is defined as the radius of the smallest sphere that contains it. Intuitively, candidates in dense regions fall within smaller spheres, while those near sparse regions yield higher rarity scores. We then rank candidates by their rarity scores and select high-rarity ones as the preliminary sparse condition set $\tilde{C}_{k}$.
Figure 3: Sparse-Aware Semantic Sampling. Candidate sparse features (blue) are optimized in the semantic space using two potentials: (a) a sparsity potential that repels candidates from reference features (orange) toward sparse regions; and (b) a diversity potential that enforces mutual repulsion among candidates to promote dispersion.
@@
Formally, for each selected candidate feature $\tilde{f}$, we define its local neighborhood as the reference samples within an $\epsilon$-radius in the semantic space, where the candidate serves as the center and neighbors are same-class reference samples. To enable gradient-based optimization, the hard neighbor count is replaced with a sigmoid-based smooth approximation $\sigma$:
@@
By minimizing this potential through gradient-based updates to the candidate feature $\tilde{f}$, the optimized features are adaptively repelled from high-density clusters and pushed into sparse regions of the semantic space.
@@
Minimizing $\Phi_{diversity}$ encourages candidate samples to spread apart within the semantic space, discouraging trivial duplication and enforcing local intra-class diversity.
@@
Before feeding the optimized sparse features into the diffusion model, we filter them based on confidence scores from the task-aware classifier to ensure the generation quality, retaining only those above a predefined lower threshold. To maintain balanced class coverage, low-confidence samples are replaced with high-fidelity GMSS samples drawn from $\mathcal{N}(\mu_{k},0.5\Sigma_{k})$. This yields a final condition set that combines sparsity, diversity, and faithfulness for more effective diffusion-based augmentation. To ensure efficiency and consistency, all condition features are synthesized offline, organized by class, before initiating the reverse diffusion process. The overall algorithm is provided in Algorithm 1 in the Supplementary Material. [^1] Supplementary materials: https://github.com/Lchenuu/SASG-DA.
@@ 5.1 Datasets
To evaluate the effectiveness of the proposed approach, we conduct data augmentation experiments on three public sEMG benchmark datasets, namely Ninapro DB2 [29], DB4 [30], and DB7 [31], as summarized in Table 1. These datasets encompass a diverse range of gesture classes and subjects, providing a comprehensive testbed for evaluating generalization. During preprocessing, the raw sEMG signals of each subject are segmented via a sliding window of length 200 ms with a step size of 40 ms, and then standardized channel-wise. To ensure consistency with previous studies [13, 12] and focus on gesture-related categories, the rest class is excluded during preprocessing. Following established protocols [13, 12] for gesture recognition based on sEMG, we adopt a cross-trial paradigm in which the training and testing sets are divided based on the trial information, as mentioned in Table 1.
Table 2: Performance comparison with SOTAs in terms of ACC (%), Pre(%), Rec(%) and F1(%) on dataset DB7 using three different backbones. The results report the mean and standard deviation across subjects. Best performances are highlighted in bold.
@@
Baseline $77.97$ p m 4.07∗∗∗ $78.86$ p m 3.63 $78.37$ p m 3.78 $78.61$ p m 3.74 $73.54$ p m 5.28∗∗∗ $74.99$ p m 3.78 $73.96$ p m 4.76 $74.46$ p m 4.55 $73.61$ p m 5.50∗∗∗ $75.10$ p m 4.67 $74.17$ p m 4.90 $74.63$ p m 4.86
Jittering & Scaling $78.05$ p m 4.82∗∗∗ $79.12$ p m 4.30 $78.44$ p m 4.45 $78.78$ p m 4.40 $74.10$ p m 5.17∗∗∗ $75.44$ p m 3.94 $74.53$ p m 4.60 $74.92$ p m 4.44 $72.28$ p m 5.87∗∗∗ $73.98$ p m 4.85 $72.88$ p m 5.24 $73.43$ p m 5.15
[…]
FreqMask $77.99$ p m 4.90∗∗∗ $79.02$ p m 4.32 $78.62$ p m 4.38 $78.82$ p m 4.47 $75.59$ p m 5.85∗∗∗ $77.11$ p m 4.04 $76.10$ p m 5.38 $76.60$ p m 5.34 $78.87$ p m 4.43∗∗∗ $80.06$ p m 3.76 $79.47$ p m 3.87 $79.76$ p m 3.97
RIM $79.04$ p m 3.97∗∗∗ $79.75$ p m 3.58 $79.49$ p m 3.66 $79.62$ p m 3.67 $76.12$ p m 5.56∗∗∗ $77.28$ p m 4.17 $76.55$ p m 5.12 $76.91$ p m 5.04 $76.93$ p m 5.25∗∗∗ $78.18$ p m 4.63 $77.48$ p m 4.71 $77.83$ p m 4.72
Dominant Shuffle $79.41$ p m 4.47∗∗∗ $80.22$ p m 3.93 $79.80$ p m 4.13 $80.01$ p m 4.08 $75.21$ p m 5.90∗∗∗ $77.03$ p m 4.15 $75.67$ p m 5.59 $76.34$ p m 5.46 $78.92$ p m 4.46∗∗∗ $80.37$ p m 3.74 $79.46$ p m 4.00 $79.91$ p m 4.00
Mixup $76.99$ p m 5.01∗∗∗ $77.79$ p m 4.55 $77.58$ p m 4.53 $77.68$ p m 4.59 $75.17$ p m 4.33∗∗∗ $76.03$ p m 3.88 $75.68$ p m 3.96 $75.85$ p m 3.94 $73.90$ p m 5.72∗∗∗ $75.18$ p m 5.02 $74.55$ p m 5.20 $74.86$ p m 5.17
STAug $77.43$ p m 5.15∗∗∗ $78.30$ p m 4.60 $77.95$ p m 4.66 $78.12$ p m 4.67 $75.32$ p m 4.32∗∗∗ $76.29$ p m 3.69 $75.72$ p m 3.99 $76.00$ p m 3.91 $73.53$ p m 5.65∗∗∗ $74.77$ p m 5.03 $74.22$ p m 5.01 $74.49$ p m 5.06
[…]
SimPSI $78.23$ p m 4.31∗∗∗ $78.94$ p m 3.89 $78.60$ p m 4.00 $78.77$ p m 3.98 $74.12$ p m 5.33∗∗∗ $75.43$ p m 3.81 $74.52$ p m 4.78 $74.97$ p m 4.69 $74.51$ p m 5.36∗∗∗ $75.92$ p m 4.58 $75.16$ p m 4.77 $75.54$ p m 4.70
E-TRGAN $78.35$ p m 4.72∗∗∗ $78.95$ p m 3.89 $78.83$ p m 4.56 $78.89$ p m 4.03 $74.16$ p m 3.77∗∗∗ $76.02$ p m 4.91 $75.85$ p m 3.68 $75.93$ p m 4.44 $74.24$ p m 3.55∗∗∗ $75.30$ p m 4.21 $74.95$ p m 3.12 $75.12$ p m 4.87
DiffMix $78.14$ p m 4.67∗∗∗ $78.90$ p m 4.15 $78.52$ p m 4.31 $78.71$ p m 4.26 $74.44$ p m 4.74∗∗∗ $75.40$ p m 4.01 $74.86$ p m 4.30 $75.13$ p m 4.19 $74.26$ p m 5.31∗∗∗ $75.51$ p m 4.60 $74.89$ p m 4.69 $75.20$ p m 4.69
SSE-LDM $78.13$ p m 4.61∗∗∗ $78.86$ p m 4.11 $78.53$ p m 4.23 $78.70$ p m 4.19 $74.64$ p m 4.54∗∗∗ $75.55$ p m 4.05 $75.03$ p m 4.20 $75.29$ p m 4.16 $73.69$ p m 5.53∗∗∗ $74.94$ p m 4.96 $74.24$ p m 5.06 $74.59$ p m 5.05
EEG-Enhancing $78.40$ p m 4.36∗∗∗ $79.27$ p m 3.87 $78.80$ p m 4.06 $79.04$ p m 3.96 $73.84$ p m 6.87∗∗∗ $75.41$ p m 4.65 $74.23$ p m 6.35 $74.79$ p m 5.54 $73.61$ p m 5.41∗∗∗ $75.10$ p m 4.47 $74.24$ p m 4.73 $74.67$ p m 4.60
NTD $78.71$ p m 3.59∗∗∗ $79.58$ p m 3.22 $79.08$ p m 3.31 $79.33$ p m 3.26 $75.86$ p m 4.64∗∗∗ $77.18$ p m 3.79 $76.23$ p m 4.26 $76.70$ p m 4.02 $77.82$ p m 4.93∗∗∗ $79.03$ p m 4.32 $78.30$ p m 4.52 $78.66$ p m 4.42
[…]
DESAM $79.73$ p m 4.20∗∗∗ $80.58$ p m 3.70 $80.19$ p m 3.84 $80.38$ p m 3.79 $75.88$ p m 4.81∗∗ $77.20$ p m 3.90 $76.18$ p m 4.38 $76.69$ p m 4.23 $78.64$ p m 5.09∗∗∗ $79.61$ p m 4.55 $79.04$ p m 4.70 $79.32$ p m 4.67
Minority $79.86$ p m 4.02∗∗∗ $80.30$ p m 3.69 $80.37$ p m 3.68 $80.33$ p m 3.76 $75.14$ p m 4.40∗∗∗ $75.86$ p m 4.08 $75.67$ p m 3.98 $75.76$ p m 4.15 $79.84$ p m 4.37∗∗∗ $80.53$ p m 3.89 $80.37$ p m 3.87 $80.45$ p m 3.94
[…]
  • $p<0.05$, ∗∗ $p<0.01$, ∗∗∗ $p<0.001$, $N=20$.
Table 3: Performance comparison with SOTAs in terms of ACC (%), Pre(%), Rec(%) and F1(%) on dataset DB4 using three different backbones. The results report the mean and standard deviation across subjects. Best performances are highlighted in bold.
@@
Baseline $65.78$ p m 5.88 $66.97$ p m 5.95 $67.49$ p m 6.15 $67.23$ p m 6.12 $62.62$ p m 4.97 $64.18$ p m 5.34 $64.34$ p m 5.18 $64.26$ p m 5.31 $61.25$ p m 7.03 $63.43$ p m 7.04 $63.05$ p m 7.43 $63.24$ p m 7.39
Jittering & Scaling $66.01$ p m 5.75 $67.35$ p m 5.85 $67.80$ p m 5.98 $67.57$ p m 5.99 $63.05$ p m 5.02 $64.60$ p m 5.28 $64.75$ p m 5.20 $64.67$ p m 5.30 $59.97$ p m 6.89 $62.23$ p m 7.01 $61.74$ p m 7.22 $61.98$ p m 7.22
[…]
FreqMask $66.67$ p m 6.16 $68.08$ p m 6.22 $68.35$ p m 6.23 $68.21$ p m 6.37 $64.50$ p m 5.55 $66.31$ p m 5.76 $66.23$ p m 5.61 $66.27$ p m 5.80 $67.08$ p m 6.25 $68.99$ p m 6.29 $68.82$ p m 6.33 $68.90$ p m 6.51
RIM $67.18$ p m 5.88 $68.31$ p m 6.06 $68.90$ p m 6.12 $68.60$ p m 6.19 $64.47$ p m 5.19 $65.75$ p m 5.39 $66.21$ p m 5.38 $65.98$ p m 5.48 $65.08$ p m 6.44 $66.58$ p m 6.33 $66.77$ p m 6.67 $66.67$ p m 6.66
Dominant Shuffle $67.21$ p m 5.77 $68.41$ p m 5.84 $68.96$ p m 5.98 $68.68$ p m 6.00 $64.22$ p m 5.59 $66.01$ p m 5.63 $66.06$ p m 5.70 $66.03$ p m 5.84 $65.89$ p m 6.58 $67.89$ p m 6.34 $67.69$ p m 6.74 $67.79$ p m 6.80
Mixup $66.21$ p m 5.68 $67.48$ p m 5.67 $67.86$ p m 5.81 $67.67$ p m 5.88 $64.50$ p m 4.81 $65.96$ p m 5.04 $66.20$ p m 5.00 $66.08$ p m 5.14 $62.13$ p m 6.84 $63.95$ p m 6.94 $63.82$ p m 7.06 $63.88$ p m 7.14
STAug $66.41$ p m 5.89 $67.77$ p m 6.03 $68.09$ p m 6.06 $67.93$ p m 6.14 $63.96$ p m 5.70 $65.31$ p m 5.88 $65.67$ p m 5.82 $65.49$ p m 5.95 $61.46$ p m 6.95 $63.36$ p m 7.12 $63.10$ p m 7.25 $63.23$ p m 7.20
[…]
SimPSI $66.72$ p m 6.02 $67.93$ p m 5.99 $68.40$ p m 6.10 $68.16$ p m 6.19 $64.06$ p m 5.43 $65.54$ p m 5.58 $65.78$ p m 5.58 $65.66$ p m 5.73 $62.44$ p m 7.51 $64.48$ p m 7.38 $64.06$ p m 7.85 $64.27$ p m 7.79
E-TRGAN $65.93$ p m 5.42 $67.70$ p m 6.11 $67.28$ p m 5.67 $67.51$ p m 5.94 $63.11$ p m 5.33 $64.60$ p m 6.48 $64.55$ p m 5.58 $64.57$ p m 6.23 $62.66$ p m 5.21 $64.50$ p m 5.89 $64.54$ p m 5.07 $64.52$ p m 6.37
DiffMix $66.08$ p m 6.07 $67.21$ p m 6.09 $67.81$ p m 6.29 $67.51$ p m 6.31 $62.14$ p m 5.64 $63.35$ p m 5.88 $63.83$ p m 5.89 $63.59$ p m 5.96 $63.22$ p m 6.45 $64.89$ p m 6.37 $64.99$ p m 6.85 $64.94$ p m 6.74
SSE-LDM $66.10$ p m 5.33 $67.16$ p m 5.39 $67.76$ p m 5.53 $67.46$ p m 5.56 $63.53$ p m 4.97 $64.76$ p m 5.32 $65.26$ p m 5.18 $65.01$ p m 5.29 $62.56$ p m 6.50 $64.18$ p m 6.48 $64.34$ p m 6.84 $64.26$ p m 6.78
EEG-Enhancing $66.48$ p m 5.84∗∗ $67.66$ p m 5.90 $68.22$ p m 6.06 $67.94$ p m 5.97 $62.71$ p m 4.95∗∗ $64.16$ p m 5.26 $64.47$ p m 5.13 $64.31$ p m 5.19 $61.86$ p m 6.80∗∗ $63.69$ p m 6.85 $63.63$ p m 7.11 $63.66$ p m 6.97
NTD $67.90$ p m 6.13∗∗ $68.92$ p m 6.18 $69.55$ p m 6.36 $69.23$ p m 6.27 $64.29$ p m 5.89∗∗ $65.72$ p m 6.21 $65.85$ p m 5.93 $65.78$ p m 6.06 $65.69$ p m 8.56∗∗ $67.19$ p m 8.31 $67.39$ p m 8.53 $67.29$ p m 8.41
PatchEMG $67.50$ p m 5.09∗∗ $68.93$ p m 5.41 $69.26$ p m 5.49 $69.09$ p m 5.44 $64.68$ p m 4.85∗∗ $66.38$ p m 5.24 $66.37$ p m 5.02 $66.37$ p m 5.14 $67.45$ p m 6.13∗∗ $69.42$ p m 6.09 $69.21$ p m 6.37 $69.31$ p m 6.33
DESAM $67.63$ p m 5.85∗∗ $68.82$ p m 5.97 $69.33$ p m 6.15 $69.07$ p m 6.16 $63.91$ p m 5.42∗∗ $65.24$ p m 5.68 $65.60$ p m 5.56 $65.42$ p m 5.69 $64.06$ p m 7.32∗∗ $65.67$ p m 7.01 $65.90$ p m 7.53 $65.78$ p m 7.41
Minority $68.39$ p m 5.94 $69.49$ p m 5.92 $70.03$ p m 6.18 $69.76$ p m 6.17 $64.47$ p m 5.78 $65.86$ p m 5.85 $66.13$ p m 5.97 $65.99$ p m 6.00 $66.74$ p m 7.20 $68.38$ p m 7.21 $68.42$ p m 7.45 $66.40$ p m 7.56
[…]
  • $p<0.05$, ∗∗ $p<0.01$, ∗∗∗ $p<0.001$, $N=10$.
Table 4: Performance comparison with SOTAs in terms of ACC (%), Pre(%), Rec(%) and F1(%) on dataset DB2 using three different backbones. The results report the mean and standard deviation across subjects. Best performances are highlighted in bold.